OpenStudy (emily_98):

In simplest radical form, what are the solutions to the quadratic equation 6 = x2 – 10x? Quadratic formula: x =

OpenStudy (anonymous):

\[x = 5 \pm \sqrt{31}\]

OpenStudy (anonymous):

If you need further explanation I can do?

OpenStudy (welshfella):

you should do that direct answers are not allowed in Open Study

OpenStudy (anonymous):

Ok let me explain then

OpenStudy (emily_98):

thanks for the opinion lol

OpenStudy (anonymous):

\[x^2 - 10x - 6 = 0 \\ x^2 - 10x + 25 - 25 - 6 = 0\] I added 25 to make it in this form \[(a + b)^2\] and to make no effect we have to subtract 25 too. so \[(x - 5)^2 - 31 = 0 \\ (x - 5)^2 - (\sqrt[2]{31})^2 = 0\\ (x - 5 + \sqrt{31})(x - 5 - \sqrt{31}) = 0 \\\] so \[x = 5 \pm \sqrt{31}\]

OpenStudy (anonymous):

i used \[(a^2 - b^2) = (a - b)(a+b)\]